OpenStudy (anonymous):

Simplify the expression shown below: ln(e^3/x^4) a. 3-x^4 b. 3-4ln(x) c. 3/x^4 d. 3e/4

OpenStudy (anonymous):

ln(e^3/x^4) you can rewrite as ln(e^3)-ln(x^4) since ln have base e , ln(e^3)=3 ln(x^4) you can rewrite as 4ln(x) formulas : ln(a/b)=ln(a)-ln(b) ln(a^n)=nln(a)

OpenStudy (anonymous):

so 'B' ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

awesome thanks man

OpenStudy (anonymous):

no problem :)